# A Survey of Reinforcement Learning For Economics

### 15 Mar 2026 v2

@@ 10.3.2 Neural Fictitious Self-Play
NFSP (Heinrich and Silver, 2016) combines fictitious play161161161Fictitious play (Brown, 1951) is a classical learning rule where each player best-responds to the empirical distribution of opponents’ past actions. Under certain conditions (e.g., two-player zero-sum games), the time-average strategies converge to Nash equilibrium. (Brown, 1951) with deep Q-learning.162162162Deep Q-Network (DQN) (Mnih et al., 2015) approximates the Q-function with a neural network, using experience replay and a target network to stabilize training. The target network $\theta^{-}$ in the loss function is a delayed copy of the main network, updated periodically. Each player maintains two networks: a best-response network $Q_{\theta}$ trained via DQN, and an average strategy network $\bar{\pi}_{\phi}$ trained via supervised learning on the best-response actions:
@@ 10.3.3 Poker Results
These methods achieved superhuman performance in poker. Deep CFR attained exploitability of 37 mbb/g163163163Exploitability measures how far a strategy is from Nash equilibrium, defined as the maximum expected gain an adversary could achieve by best-responding. Zero exploitability means the strategy is unexploitable (Nash). In poker, exploitability is measured in milli-big-blinds per game (mbb/g). in heads-up flop hold’em (Brown et al., 2019). Libratus defeated top human professionals in heads-up no-limit hold’em using nested subgame solving with CFR (Brown and Sandholm, 2018). Pluribus extended to six-player no-limit hold’em (Brown and Sandholm, 2019). The game tree of heads-up no-limit hold’em contains $\sim 10^{161}$ states; these results demonstrate that CFR-based methods scale to economically relevant game sizes.
@@ 10.4 The Coase Conjecture
Coase (1972) conjectured that a durable goods monopolist164164164A durable good provides utility over multiple periods (e.g., a car, appliance, or software license) rather than being consumed immediately. Unlike non-durable goods, durable goods create intertemporal competition, as the seller at time $t$ competes with her own future self at $t+1$. loses market power when buyers are patient. Unable to commit to future prices, the seller competes with her future self, eroding rents. As the inter-offer interval shrinks ($\delta\to 1$), price collapses to marginal cost and all surplus is extracted by buyers. Gul et al. (1986) formalized this for the gap case, where the seller’s cost is strictly below all buyer valuations, guaranteeing trade and a unique stationary equilibrium.165165165Stokey (1981) showed that a monopolist facing rational consumers cannot price discriminate intertemporally; Bulow (1982) proved that in the no-gap case, the monopolist prefers renting to selling. The no-gap case admits multiple equilibria with more complex dynamics.
@@ 10.4.2 Equilibrium Analysis
For $\pi<\pi^{*}$, the seller pools (offers $v_{L}$ immediately). For $\pi>\pi^{*}$, the seller screens (offers $P^{*}$, then $v_{L}$ if rejected).166166166In a screening equilibrium, the seller uses price to separate buyer types: high-value buyers accept immediately at a high price, while low-value buyers reject and receive a lower offer. In a pooling equilibrium, the seller offers a single price that all types accept. The gap case ($0=c<v_{L}$) guarantees trade in equilibrium. The Coase conjecture manifests as $\delta\to 1$, where $P^{*}(\delta)\to v_{L}$ and the seller cannot extract surplus from high types.
@@ 10.4.3 Computational Results
I model the bargaining game as an extensive-form game and apply CFR.167167167The game tree has two information sets for the seller (indexed by rejection history) and two for the buyer (indexed by private type). CFR runs for 5,000 iterations per parameter configuration; NashConv (sum of exploitabilities) measures convergence.
@@ 11 Bandits and Dynamic Pricing
This chapter focuses on online reinforcement learning (Section 2), where exploration is a real cost and the agent must balance exploration against exploitation. A seller faces $T$ customers in sequence, setting a price for each and observing only whether the customer bought.169169169Rothschild (1974) posed pricing under demand uncertainty as a two-armed bandit problem. His insight was that a myopic seller can get stuck at a suboptimal price forever, because exploiting the currently best-looking price generates no information about alternatives. The seller does not observe the customer’s willingness to pay. Regret, the total revenue gap between the seller’s policy and the best fixed price in hindsight, is the central measure: if the optimal fixed price earns $r^{*}$ per customer, a policy with regret $R(T)$ earns $Tr^{*}-R(T)$ in total. The central question is how fast $R(T)$ shrinks as the seller accumulates purchase data, and how structural assumptions about demand affect this rate.
@@ 11.1.1 No Structure on Demand
Kleinberg and Leighton (2003) study the simplest version of the problem. Customers arrive with valuations drawn i.i.d. from an unknown distribution on $[0,1]$, and the seller posts a price from a continuous set. The demand curve $D(p)=\Pr(v\geq p)$ is unknown; the only feedback is whether each customer bought. The seller’s goal is to minimize regret against the single price $p^{*}$ that maximizes $p\cdot D(p)$. Kleinberg and Leighton prove that the minimax regret is $\Theta(\sqrt{T})$.170170170The upper bound, $O(\sqrt{T\log T})$, discretizes $[0,1]$ into $K=\lceil(T/\log T)^{1/4}\rceil$ prices and runs the UCB1 algorithm of Auer et al. (2002a). The lower bound, $\Omega(\sqrt{T})$, constructs a family of demand curves parameterized by the location of the optimal price $p^{*}\in[0.3,0.4]$. The key tension: posting prices far from $p^{*}$ is informative about demand but costly in revenue; posting prices near $p^{*}$ is cheap but uninformative. Resolving this tension costs at least $\Omega(\sqrt{T})$ in cumulative revenue. UCB1 (Auer et al., 2002a) selects the price maximizing $\hat{\mu}_{p_{k}}(t)+\sqrt{2\ln t/N_{p_{k}}(t)}$, where $\hat{\mu}_{p_{k}}(t)$ is the empirical mean profit and $N_{p_{k}}(t)$ the number of trials; the second term is an exploration bonus that shrinks as a price is tried more, implementing the principle of optimism in the face of uncertainty. In concrete terms, after 10,000 customers the seller loses roughly 100 customers’ worth of revenue to the uncertainty in demand. No algorithm can do better without imposing structure on $D$. For adversarial valuations (chosen by a worst-case opponent rather than drawn from a fixed distribution), the minimax regret rises to $\Theta(T^{2/3})$, achieved by the Exp3 algorithm171171171Exp3 (Auer et al., 2002b) maintains a weight $w_{k}(t)$ for each price, selecting $p_{k}$ with probability proportional to $w_{k}(t)$ and updating the chosen price’s weight by $\exp(\eta\hat{r}_{k,t})$ where $\hat{r}_{k,t}$ is the revenue importance-weighted by the selection probability; because no model of demand is assumed, the guarantee holds against an adversary who chooses valuations after observing the algorithm. on a discretized price grid. I focus on the stochastic setting throughout this chapter.
@@ 11.1.2 Parametric Demand
Broder and Rusmevichientong (2012) consider a parametric demand model $d(p;z)=\Pr(V\geq p)$, where $z\in\mathbb{R}^{n}$ is an unknown parameter governing the demand curve. The seller observes binary purchase decisions and updates a maximum likelihood estimate of $z$. Under standard regularity conditions (bounded demand, unique optimal price, smooth revenue function), the minimax regret is again $\Theta(\sqrt{T})$.172172172The lower bound (Theorem 3.1 of Broder and Rusmevichientong (2012)) constructs a linear demand family where all demand curves pass through the same point at the optimal price $p^{*}(z_{0})$. Observing purchases at this price provides no information about $z$. An MLE-Cycle policy that interleaves dedicated exploration rounds with greedy pricing achieves the matching upper bound $O(\sqrt{T})$ (Theorem 3.6). Parametric structure alone does not break the $\sqrt{T}$ barrier.
The picture changes under a “well-separated” condition requiring that every price in the feasible set is informative about the demand parameter. Formally, the Fisher information $I(p,z)$ is bounded below by $c_{f}>0$ for all prices $p$ and parameters zz.173173173This means no two demand curves $d(p;z)$ and $d(p;z^{\prime})$ cross on the pricing interval, so every purchase observation distinguishes the two hypotheses. In the lower-bound family of Theorem 3.1, the curves all cross at $p=1$, which is why the $\sqrt{T}$ floor emerges. Under well-separation, a pure greedy policy (MLE-Greedy, pricing at $p^{*}(\hat{z}_{t})$ each period) achieves $O(\log T)$ regret (Theorem 4.8), because every price is informative and dedicated exploration rounds become unnecessary.
@@ 11.1.3 High-Dimensional Features with Sparsity
Javanmard and Nazerzadeh (2019) extend the setting to products described by $d$-dimensional feature vectors $x_{t}$. The market value is $v_{t}=x_{t}^{\top}\theta_{0}+z_{t}$, where $\theta_{0}\in\mathbb{R}^{d}$ is unknown and $z_{t}$ is i.i.d. noise from a known log-concave distribution FF.174174174Log-concavity of $F$ and 1F1-F is satisfied by the normal, logistic, uniform, Laplace, and exponential distributions. It ensures that expected revenue $p\cdot[1-F(p-x_{t}^{\top}\theta_{0})]$ is strictly quasi-concave in $p$, giving a unique optimal price. Only $s_{0}$ of the $d$ coordinates of $\theta_{0}$ are nonzero, but the seller does not know which ones.
Their RMLP algorithm (Regularized Maximum Likelihood Pricing) operates in episodes whose lengths double (1, 2, 4, 8, …). At each episode boundary, the seller fits a LASSO-penalized maximum likelihood estimate of $\theta_{0}$ using data from the previous episode, then prices greedily throughout the current episode. The regret is $O(s_{0}\log d\cdot\log T)$ (Theorem 4.1), with a matching lower bound of $\Omega(s_{0}(\log d+\log T))$ (Theorem 5.1). The reason this beats $\sqrt{T}$ connects back to the Broder lower bound. In the parametric model of Section 11.1.2, all demand curves can cross at the optimal price, making that price uninformative. Here, customer features vary across periods, so the aggregate demand function at any fixed price changes in proportion to the estimation error in $\theta_{0}$. Every price is informative, and dedicated exploration rounds become unnecessary.175175175If some feature directions are rarely observed, the seller cannot learn all coordinates of $\theta_{0}$ quickly, and regret degrades to $O(\sqrt{\log(d)\cdot T})$ (Theorem 4.2). If the noise distribution belongs to a known parametric family but its scale parameter is unknown, regret reverts to $\Omega(\sqrt{T})$ (Theorem 7.1), foreshadowing the result of Xu and Wang (2021) discussed in Section 11.3. Even with hundreds of features, if only a handful matter, learning is fast.
@@ 11.2 Revealed Preference and Partial Identification
if $p_{k}$ is not dominated, and $I_{p_{k}}(t)=0$ otherwise, where $\hat{\mu}_{p_{k}}(t)$ is the average profit observed at price $p_{k}$ and $N_{p_{k}}(t)$ is the number of trials.176176176Standard UCB1 uses an exploration bonus of $\sqrt{2\ln t/N_{p_{k}}(t)}$, which assumes rewards in $[0,1]$. Since profit at price $p_{k}$ is bounded by $p_{k}$, scaling the bonus by $p_{k}$ tightens exploration for cheap prices that cannot contribute much profit regardless. Two innovations drive the improvement over standard UCB1. First, dominance elimination reduces the effective number of arms the algorithm must explore. Second, the $p_{k}$ scaling focuses exploration on prices where uncertainty actually matters for profit. Together, these yield $O(\log T)$ regret.
@@
where $\Delta_{k}=\mu_{k^{*}}-\mu_{k}$ is the gap between arm $k$ and the optimal arm.177177177The first sum is the leading term, scaling as $O(\log T)$. The second sum is a constant that does not grow with $T$. Replacing $p_{k}$ with 1 recovers the standard UCB1 bound, which is looser.
Misra et al. (2019) calibrate the model to a field experiment at ZipRecruiter, a B2B online recruiting platform. With 7,870 customers per month, 1,000 segments, and 10 price points from \$19 to \$399, UCB-PI achieves 98% of oracle profit and produces 43% higher profits during the first month of testing compared to a learn-then-earn alternative. The algorithm has both higher mean profit and lower variance, reflecting the value of eliminating dominated prices early.178178178For multi-product settings, Mueller et al. (2019) impose low-rank structure on the price-sensitivity matrix, achieving regret $O(T^{3/4}\sqrt{d})$ that scales with the latent demand dimension $d$ rather than the number of products. Badanidiyuru et al. (2013) extend the bandit framework to handle inventory constraints (“bandits with knapsacks”), relevant when the seller faces limited stock alongside the pricing decision.
@@ 11.3 The Value of Knowing the Noise Distribution
Xu and Wang (2021) ask how much it helps to know the shape of demand uncertainty. In their model, a feature vector $x_{t}\in\mathbb{R}^{d}$ describes each sales session, the customer’s valuation is $w_{t}=x_{t}^{\top}\theta^{*}+N_{t}$ where $\theta^{*}$ is unknown and $N_{t}$ is zero-mean i.i.d. noise with CDF $F$, and the seller observes only whether the customer bought at the posted price. The answer depends on whether $F$ is known.179179179The regret benchmark here differs from Section 11.1.1. Kleinberg and Leighton (2003) and Broder and Rusmevichientong (2012) measure regret against the best fixed price; Xu and Wang (2021) and Liu et al. (2024) measure regret against the clairvoyant contextual policy that sets the optimal price $p_{t}^{*}$ for each customer’s features $x_{t}$. The contextual benchmark is harder, which makes the $O(d\log T)$ rate more impressive.
@@
If $F$ is known (the seller knows that demand shocks are, say, normally distributed with known variance), the EMLP algorithm (Epoch-based Maximum Likelihood Pricing) achieves regret $O(d\log T)$ (Theorem 3).180180180EMLP runs in doubling epochs of length $\tau_{k}=2^{k-1}$. At each epoch boundary, the seller fits a maximum likelihood estimate $\hat{\theta}_{k}$ using data from the previous epoch, then prices greedily at $p_{t}=J(x_{t}^{\top}\hat{\theta}_{k})$ throughout the epoch, where $J(u)=\arg\max_{v}\,v[1-F(v-u)]$ is the revenue-maximizing price function. The key technical insight is that the negative log-likelihood is strongly convex (Lemma 7 of Xu and Wang (2021)), so MLE concentrates at rate $O(d/\tau_{k})$. Since regret is quadratic in the parameter estimation error (Lemma 5) and there are $O(\log T)$ epochs, the total regret is $O(d\log T)$. After 10,000 customers with $d=5$ features, the revenue loss is roughly 50 customers’ worth, an improvement over the $\sqrt{T}\approx 100$ customers’ worth that Kleinberg’s lower bound imposes without structural knowledge.
@@
If $F$ is unknown (even if only the variance of a Gaussian is unknown, with everything else known), the regret is at least $\Omega(\sqrt{T})$ (Theorem 12).181181181The lower bound constructs two noise variances $\sigma_{1}=1$ and $\sigma_{2}=1-T^{-1/4}$. Any algorithm that performs well under both must spend $\Omega(\sqrt{T})$ revenue distinguishing the two cases. This extends the “uninformative price” phenomenon of Broder and Rusmevichientong (2012): when the seller does not know $F$, there exist prices at which observed purchase behavior is nearly identical under different demand parameters. The seller is back to the Kleinberg baseline, with no algorithm able to achieve sublinear improvement regardless of how many features are available.
@@
The gap between $O(d\log T)$ and $\Omega(\sqrt{T})$ is super-polynomial in $T$, not merely a constant factor. When an economist specifies a logit or probit demand model, the assumed noise distribution purchases a qualitative improvement in learning rate. Semiparametric approaches that leave the error distribution unspecified pay a concrete cost, reverting from logarithmic to polynomial regret.182182182Tullii et al. (2024) establish the tightest known bound under minimal distributional assumptions: if the noise distribution (c.d.f.) is merely Lipschitz continuous, the minimax regret is $\Theta(T^{2/3})$, strictly between the $\log T$ rate with known $F$ and the $\sqrt{T}$ rate with unknown $F$. Fan et al. (2024) consider a semiparametric setting where the noise density is smooth and connect the pricing problem to the econometrics of semiparametric estimation.

### 17 Mar 2026 v3

@@ 9.3.2 Neural Fictitious Self-Play
NFSP (Heinrich and Silver, 2016) combines fictitious play168168168Fictitious play (Brown, 1951) is a classical learning rule where each player best-responds to the empirical distribution of opponents’ past actions. Under certain conditions (e.g., two-player zero-sum games), the time-average strategies converge to Nash equilibrium. (Brown, 1951) with deep Q-learning.169169169Deep Q-Network (DQN) (Mnih et al., 2015) approximates the Q-function with a neural network, using experience replay and a target network to stabilize training. The target network $\theta^{-}$ in the loss function is a delayed copy of the main network, updated periodically. Each player maintains two networks: a best-response network $Q_{\theta}$ trained via DQN, and an average strategy network $\bar{\pi}_{\phi}$ trained via supervised learning on the best-response actions:
@@ 9.3.3 Poker Results
These methods achieved superhuman performance in poker. Deep CFR attained exploitability of 37 mbb/g170170170Exploitability measures how far a strategy is from Nash equilibrium, defined as the maximum expected gain an adversary could achieve by best-responding. Zero exploitability means the strategy is unexploitable (Nash). In poker, exploitability is measured in milli-big-blinds per game (mbb/g). in heads-up flop hold’em (Brown et al., 2019). Libratus defeated top human professionals in heads-up no-limit hold’em using nested subgame solving with CFR (Brown and Sandholm, 2018). Pluribus extended to six-player no-limit hold’em (Brown and Sandholm, 2019). The game tree of heads-up no-limit hold’em contains $\sim 10^{161}$ states; these results demonstrate that CFR-based methods scale to economically relevant game sizes.
@@ 9.4 The Coase Conjecture
Coase (1972) conjectured that a durable goods monopolist171171171A durable good provides utility over multiple periods (e.g., a car, appliance, or software license) rather than being consumed immediately. Unlike non-durable goods, durable goods create intertemporal competition, as the seller at time $t$ competes with her own future self at $t+1$. loses market power when buyers are patient. Unable to commit to future prices, the seller competes with her future self, eroding rents. As the inter-offer interval shrinks ($\delta\to 1$), price collapses to marginal cost and all surplus is extracted by buyers. Gul et al. (1986) formalized this for the gap case, where the seller’s cost is strictly below all buyer valuations, guaranteeing trade and a unique stationary equilibrium.172172172Stokey (1981) showed that a monopolist facing rational consumers cannot price discriminate intertemporally; Bulow (1982) proved that in the no-gap case, the monopolist prefers renting to selling. The no-gap case admits multiple equilibria with more complex dynamics.
@@ 9.4.2 Equilibrium Analysis
For $\pi<\pi^{*}$, the seller pools (offers $v_{L}$ immediately). For $\pi>\pi^{*}$, the seller screens (offers $P^{*}$, then $v_{L}$ if rejected).173173173In a screening equilibrium, the seller uses price to separate buyer types: high-value buyers accept immediately at a high price, while low-value buyers reject and receive a lower offer. In a pooling equilibrium, the seller offers a single price that all types accept. The gap case ($0=c<v_{L}$) guarantees trade in equilibrium. The Coase conjecture manifests as $\delta\to 1$, where $P^{*}(\delta)\to v_{L}$ and the seller cannot extract surplus from high types.
@@ 9.4.3 Computational Results
I model the bargaining game as an extensive-form game and apply CFR.174174174The game tree has two information sets for the seller (indexed by rejection history) and two for the buyer (indexed by private type). CFR runs for 5,000 iterations per parameter configuration; NashConv (sum of exploitabilities) measures convergence.
@@ 10 Bandits and Dynamic Pricing
This chapter focuses on online reinforcement learning (Section 2), where, unlike the simulator-based training of preceding chapters, the agent learns directly from interactions with real customers and exploration is a real cost that must be balanced against exploitation. A seller faces $T$ customers in sequence, setting a price for each and observing only whether the customer bought.176176176Rothschild (1974) posed pricing under demand uncertainty as a two-armed bandit problem. His insight was that a myopic seller can get stuck at a suboptimal price forever, because exploiting the currently best-looking price generates no information about alternatives. The seller does not observe the customer’s willingness to pay. Regret, the total revenue gap between the seller’s policy and the best fixed price in hindsight, is the central measure: if the optimal fixed price earns $r^{*}$ per customer, a policy with regret $R(T)$ earns $Tr^{*}-R(T)$ in total. The central question is how fast $R(T)$ shrinks as the seller accumulates purchase data, and how structural assumptions about demand affect this rate.
@@ 10.1.1 No Structure on Demand
Kleinberg and Leighton (2003) study the simplest version of the problem. Customers arrive with valuations drawn i.i.d. from an unknown distribution on $[0,1]$, and the seller posts a price from a continuous set. The demand curve $D(p)=\Pr(v\geq p)$ is unknown; the only feedback is whether each customer bought. The seller’s goal is to minimize regret against the single price $p^{*}$ that maximizes $p\cdot D(p)$. Kleinberg and Leighton prove that the minimax regret is $\Theta(\sqrt{T})$.177177177The upper bound, $O(\sqrt{T\log T})$, discretizes $[0,1]$ into $K=\lceil(T/\log T)^{1/4}\rceil$ prices and runs the UCB1 algorithm of Auer et al. (2002a). The lower bound, $\Omega(\sqrt{T})$, constructs a family of demand curves parameterized by the location of the optimal price $p^{*}\in[0.3,0.4]$. The key tension: posting prices far from $p^{*}$ is informative about demand but costly in revenue; posting prices near $p^{*}$ is cheap but uninformative. Resolving this tension costs at least $\Omega(\sqrt{T})$ in cumulative revenue. UCB1 (Auer et al., 2002a) selects the price maximizing $\hat{\mu}_{p_{k}}(t)+\sqrt{2\ln t/N_{p_{k}}(t)}$, where $\hat{\mu}_{p_{k}}(t)$ is the empirical mean profit and $N_{p_{k}}(t)$ the number of trials; the second term is an exploration bonus that shrinks as a price is tried more, implementing the principle of optimism in the face of uncertainty. In concrete terms, after 10,000 customers the seller loses roughly 100 customers’ worth of revenue to the uncertainty in demand. No algorithm can do better without imposing structure on $D$. For adversarial valuations (chosen by a worst-case opponent rather than drawn from a fixed distribution), the minimax regret rises to $\Theta(T^{2/3})$, achieved by the Exp3 algorithm178178178Exp3 (Auer et al., 2002b) maintains a weight $w_{k}(t)$ for each price, selecting $p_{k}$ with probability proportional to $w_{k}(t)$ and updating the chosen price’s weight by $\exp(\eta\hat{r}_{k,t})$ where $\hat{r}_{k,t}$ is the revenue importance-weighted by the selection probability; because no model of demand is assumed, the guarantee holds against an adversary who chooses valuations after observing the algorithm. on a discretized price grid. I focus on the stochastic setting throughout this chapter.
@@ 10.1.2 Parametric Demand
Broder and Rusmevichientong (2012) consider a parametric demand model $d(p;z)=\Pr(V\geq p)$, where $z\in\mathbb{R}^{n}$ is an unknown parameter governing the demand curve. The seller observes binary purchase decisions and updates a maximum likelihood estimate of $z$. Under standard regularity conditions (bounded demand, unique optimal price, smooth revenue function), the minimax regret is again $\Theta(\sqrt{T})$.179179179The lower bound (Theorem 3.1 of Broder and Rusmevichientong (2012)) constructs a linear demand family where all demand curves pass through the same point at the optimal price $p^{*}(z_{0})$. Observing purchases at this price provides no information about $z$. An MLE-Cycle policy that interleaves dedicated exploration rounds with greedy pricing achieves the matching upper bound $O(\sqrt{T})$ (Theorem 3.6). Parametric structure alone does not break the $\sqrt{T}$ barrier.
The picture changes under a “well-separated” condition requiring that every price in the feasible set is informative about the demand parameter. Formally, the Fisher information $I(p,z)$ is bounded below by $c_{f}>0$ for all prices $p$ and parameters zz.180180180This means no two demand curves $d(p;z)$ and $d(p;z^{\prime})$ cross on the pricing interval, so every purchase observation distinguishes the two hypotheses. In the lower-bound family of Theorem 3.1, the curves all cross at $p=1$, which is why the $\sqrt{T}$ floor emerges. Under well-separation, a pure greedy policy (MLE-Greedy, pricing at $p^{*}(\hat{z}_{t})$ each period) achieves $O(\log T)$ regret (Theorem 4.8), because every price is informative and dedicated exploration rounds become unnecessary.
@@ 10.1.3 High-Dimensional Features with Sparsity
Javanmard and Nazerzadeh (2019) extend the setting to products described by $d$-dimensional feature vectors $x_{t}$. The market value is $v_{t}=x_{t}^{\top}\theta_{0}+z_{t}$, where $\theta_{0}\in\mathbb{R}^{d}$ is unknown and $z_{t}$ is i.i.d. noise from a known log-concave distribution FF.181181181Log-concavity of $F$ and 1F1-F is satisfied by the normal, logistic, uniform, Laplace, and exponential distributions. It ensures that expected revenue $p\cdot[1-F(p-x_{t}^{\top}\theta_{0})]$ is strictly quasi-concave in $p$, giving a unique optimal price. Only $s_{0}$ of the $d$ coordinates of $\theta_{0}$ are nonzero, but the seller does not know which ones.
Their RMLP algorithm (Regularized Maximum Likelihood Pricing) operates in episodes whose lengths double (1, 2, 4, 8, …). At each episode boundary, the seller fits a LASSO-penalized maximum likelihood estimate of $\theta_{0}$ using data from the previous episode, then prices greedily throughout the current episode. The regret is $O(s_{0}\log d\cdot\log T)$ (Theorem 4.1), with a matching lower bound of $\Omega(s_{0}(\log d+\log T))$ (Theorem 5.1). The reason this beats $\sqrt{T}$ connects back to the Broder lower bound. In the parametric model of Section 10.1.2, all demand curves can cross at the optimal price, making that price uninformative. Here, customer features vary across periods, so the aggregate demand function at any fixed price changes in proportion to the estimation error in $\theta_{0}$. Every price is informative, and dedicated exploration rounds become unnecessary.182182182If some feature directions are rarely observed, the seller cannot learn all coordinates of $\theta_{0}$ quickly, and regret degrades to $O(\sqrt{\log(d)\cdot T})$ (Theorem 4.2). If the noise distribution belongs to a known parametric family but its scale parameter is unknown, regret reverts to $\Omega(\sqrt{T})$ (Theorem 7.1), foreshadowing the result of Xu and Wang (2021) discussed in Section 10.3. Even with hundreds of features, if only a handful matter, learning is fast.
@@ 10.2 Revealed Preference and Partial Identification
if $p_{k}$ is not dominated, and $I_{p_{k}}(t)=0$ otherwise, where $\hat{\mu}_{p_{k}}(t)$ is the average profit observed at price $p_{k}$ and $N_{p_{k}}(t)$ is the number of trials.183183183Standard UCB1 uses an exploration bonus of $\sqrt{2\ln t/N_{p_{k}}(t)}$, which assumes rewards in $[0,1]$. Since profit at price $p_{k}$ is bounded by $p_{k}$, scaling the bonus by $p_{k}$ tightens exploration for cheap prices that cannot contribute much profit regardless. Two innovations drive the improvement over standard UCB1. First, dominance elimination reduces the effective number of arms the algorithm must explore. Second, the $p_{k}$ scaling focuses exploration on prices where uncertainty actually matters for profit. Together, these yield $O(\log T)$ regret.
@@
where $\Delta_{k}=\mu_{k^{*}}-\mu_{k}$ is the gap between arm $k$ and the optimal arm.184184184The first sum is the leading term, scaling as $O(\log T)$. The second sum is a constant that does not grow with $T$. Replacing $p_{k}$ with 1 recovers the standard UCB1 bound, which is looser.
Misra et al. (2019) calibrate the model to a field experiment at ZipRecruiter, a B2B online recruiting platform. With 7,870 customers per month, 1,000 segments, and 10 price points from \$19 to \$399, UCB-PI achieves 98% of oracle profit and produces 43% higher profits during the first month of testing compared to a learn-then-earn alternative. The algorithm has both higher mean profit and lower variance, reflecting the value of eliminating dominated prices early.185185185For multi-product settings, Mueller et al. (2019) impose low-rank structure on the price-sensitivity matrix, achieving regret $O(T^{3/4}\sqrt{d})$ that scales with the latent demand dimension $d$ rather than the number of products. Badanidiyuru et al. (2013) extend the bandit framework to handle inventory constraints (“bandits with knapsacks”), relevant when the seller faces limited stock alongside the pricing decision.
@@ 10.3 The Value of Knowing the Noise Distribution
Xu and Wang (2021) ask how much it helps to know the shape of demand uncertainty. In their model, a feature vector $x_{t}\in\mathbb{R}^{d}$ describes each sales session, the customer’s valuation is $w_{t}=x_{t}^{\top}\theta^{*}+N_{t}$ where $\theta^{*}$ is unknown and $N_{t}$ is zero-mean i.i.d. noise with CDF $F$, and the seller observes only whether the customer bought at the posted price. The answer depends on whether $F$ is known.186186186The regret benchmark here differs from Section 10.1.1. Kleinberg and Leighton (2003) and Broder and Rusmevichientong (2012) measure regret against the best fixed price; Xu and Wang (2021) and Liu et al. (2024) measure regret against the clairvoyant contextual policy that sets the optimal price $p_{t}^{*}$ for each customer’s features $x_{t}$. The contextual benchmark is harder, which makes the $O(d\log T)$ rate more impressive.
@@
If $F$ is known (the seller knows that demand shocks are, say, normally distributed with known variance), the EMLP algorithm (Epoch-based Maximum Likelihood Pricing) achieves regret $O(d\log T)$ (Theorem 3).187187187EMLP runs in doubling epochs of length $\tau_{k}=2^{k-1}$. At each epoch boundary, the seller fits a maximum likelihood estimate $\hat{\theta}_{k}$ using data from the previous epoch, then prices greedily at $p_{t}=J(x_{t}^{\top}\hat{\theta}_{k})$ throughout the epoch, where $J(u)=\arg\max_{v}\,v[1-F(v-u)]$ is the revenue-maximizing price function. The key technical insight is that the negative log-likelihood is strongly convex (Lemma 7 of Xu and Wang (2021)), so MLE concentrates at rate $O(d/\tau_{k})$. Since regret is quadratic in the parameter estimation error (Lemma 5) and there are $O(\log T)$ epochs, the total regret is $O(d\log T)$. After 10,000 customers with $d=5$ features, the revenue loss is roughly 50 customers’ worth, an improvement over the $\sqrt{T}\approx 100$ customers’ worth that Kleinberg’s lower bound imposes without structural knowledge.
@@
If $F$ is unknown (even if only the variance of a Gaussian is unknown, with everything else known), the regret is at least $\Omega(\sqrt{T})$ (Theorem 12).188188188The lower bound constructs two noise variances $\sigma_{1}=1$ and $\sigma_{2}=1-T^{-1/4}$. Any algorithm that performs well under both must spend $\Omega(\sqrt{T})$ revenue distinguishing the two cases. This extends the “uninformative price” phenomenon of Broder and Rusmevichientong (2012): when the seller does not know $F$, there exist prices at which observed purchase behavior is nearly identical under different demand parameters. The seller is back to the Kleinberg baseline, with no algorithm able to achieve sublinear improvement regardless of how many features are available.
@@
The gap between $O(d\log T)$ and $\Omega(\sqrt{T})$ is super-polynomial in $T$, not merely a constant factor. When an economist specifies a logit or probit demand model, the assumed noise distribution purchases a qualitative improvement in learning rate. Semiparametric approaches that leave the error distribution unspecified pay a concrete cost, reverting from logarithmic to polynomial regret.189189189Tullii et al. (2024) establish the tightest known bound under minimal distributional assumptions: if the noise distribution (c.d.f.) is merely Lipschitz continuous, the minimax regret is $\Theta(T^{2/3})$, strictly between the $\log T$ rate with known $F$ and the $\sqrt{T}$ rate with unknown $F$. Fan et al. (2024) consider a semiparametric setting where the noise density is smooth and connect the pricing problem to the econometrics of semiparametric estimation.